OpenStudy (anonymous):

1/7(B+35)3 linear equations

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

can you help me @RunawayGalaxy

OpenStudy (anonymous):

Well you're not being too clear in the question you're asking here, do you want to solve for B ?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

I don't understand what the problem is. There is no equal sign, so this is not an equation. what do you need to do?

OpenStudy (anonymous):

ohh so sorry =3

OpenStudy (mathstudent55):

Ok. Let me write it again.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Solve for B: \(\large \dfrac{1}{7}(B + 35) = 3\)

OpenStudy (mathstudent55):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Ok. You have a multiplication of 1/7 on the left side that we need to get rid of. To get rid of a multiplication, you multiply by its reciprocal. The reciprocal of 1/7 is 7, so we multiply both sides by 7. Can you do that and show what you get?

OpenStudy (mathstudent55):

\(\large 7 \times \dfrac{1}{7}(B + 35) = 7 \times 3\) What do you get?

OpenStudy (anonymous):

7b and 245

OpenStudy (mathstudent55):

No. On the let side, what is 7 * 1/7? The whole idea of multiplying both sides by 7 is to get rid of 1/7 on the left side. What is 7 * 1/7?

OpenStudy (anonymous):

1

OpenStudy (mathstudent55):